Hiking up a mountain, you notice that the air temperature drops 10 degrees C for every 1,000 meters increase in elevation. Write a multiplication expression to represent the decrease in temperature if you hike up the mountain 3,000 meters. Then evaluate the expression and explain its meaning.

Thank you in advance for any help you can provide.

Answer by stanbon(75887)

Hiking up a mountain, you notice that the air temperature drops 10 degrees C for every 1,000 meters increase in elevation. Write a multiplication expression to represent the decrease in temperature if you hike up the mountain 3,000 meters. Then evaluate the expression and explain its meaning.
-------------------
Let Tb be "temperature at the base of the mountain".
Let T(a) be "temperature at some altitude above the base".
Let a be the altitute above the base
------------
T(a) = T(b) - (a/1000)10
T(3000) = T(b) - (3000/1000)10
T(3000) = T(b) - 30 degrees
=================================
Cheers,
Stan H.